# spacing=nonfrench

### v2

@@ 6.1 Evaluations on Accuracy
Methodology. We first study whether SpecFS can accurately generate code to achieve self-evolution, that is, whether the code generation can accurately conform to the logic described by the specification and accurately meet the functionality of the corresponding file system module. Our test cases comprise all the modules to implement the complete logic of AtomFS [zou2019atomfs]. Specifically, to evaluate the accuracy, we first define 45 distinct modules within AtomFS and manually author their ground-truth implementations. A generated module is considered correct if it passes all functional tests for AtomFS and is deemed logically equivalent to the ground-truth through manual inspection. We implement two versions of baselines based on a few-shot learning approach. The normal version incorporates a description of the file correspondence logic and the APIs of the dependency modules; The oracle version not only includes the dependency module’s APIs but also integrates the ground-truth code of these modules as part of the prompt. We test generation accuracy using four LLMs of decreasing capability: Gemini-2.5-Pro [gemini], DeepSeek-V3.1 Reasoning [deepseekai2025deepseekv3technicalreport], GPT-5-minimal [gpt5], and Qwen3-32B [yang2025qwen3technicalreport], whose performance is ranked according to the LiveCodeBench leaderboard [livecodebench-leaderboard].
@@ 6.2 Evaluations on Generalizability
Generalizability for various locking methods. We evaluate whether the concurrency specifications of SysSpec are general for various locking methods. To this end, we utilize the dentry_lookup operation in the VFS layer of Linux, as it exhibits locking requirements at multiple granularities: a lock must be acquired for the entire hash list during traversal, while individual locks are also required for each dentry upon access. In our concurrency specification, we explicitly designate the use of two distinct locking mechanisms—lock-free RCU for the hash list and spinlocks for individual dentries. Our experiments verify that the generated code correctly adheres to the specified concurrency semantics, successfully producing code that implements multi-granularity and multi-method locking logic as intended. See details about the specification and the generated code in the Appendix.
@@ 6.6 Limitations and Discussion
Missing FS features. Although SpecFS represents a successful endeavor in realizing a complex file system using SysSpec, SpecFS is currently implemented as a user-space file system based on FUSE [fuse]. Consequently, it does not operate in kernel mode and lacks a storage stack, such as direct disk access, nor does it consider crash consistency. In the future, we will attempt to apply SpecFS to the self-evolution of industrial file systems (e.g., EROFS [gao2019erofs] or Ext4 [linux-ext4]).
@@ 7 Related Work
Other research endeavors focus on LLM-based code generation for specific domains outside of file systems. E.g., QiMeng Xpiler [dong2025qimeng] utilizes LLMs to construct a transcompiler that adapts low-level tensor programs for various hardware platforms. Autoverus [yang2025autoverus] employs LLMs to build a tool for proving the correctness of Rust code. OSVBench [li2025osvbenchbenchmarkingllmsspecification] and SpecGen [ma2025specgenautomatedgenerationformal] focus on generate formal specifications for a given code snippet.
Formal verification methods. There exists a significant body of work on formally verifying complex systems, such as operating systems [gerwin2009sel4], file systems [zou2019atomfs, zou2024reffs, chen2016fscq], or cloud systems [sun2024anvil]. These efforts also involve formal specifications to ascertain the correctness of handcrafted implementations, but cannot be used diretly for generative file systems. We can draw upon these existing specifications when authoring SysSpec for code generation.
@@ 8 Conclusion
This paper presents SysSpec and SpecFS, a framework and a case for generative file systems. Although different from traditional paradigms, SysSpec shifts the developer’s focus from writing low-level code to designing high-level specification, and SpecFS shows the potential benefits on evolvability.
@@ A.1 Specification
struct hlist_head { /* ... */ };
struct hlist_node { /* ... */ };
@@ A.2 Code Snippet
Following is the code generated by Qwen3-32B model.
@@
struct hlist_head *bucket = d_hash(parent, hash);
@@
struct dentry *dentry;
@@
while (node) {
dentry = hlist_entry(node, struct dentry, d_hash);
@@
spin_lock(&dentry->d_lock);
@@
if (dentry->d_parent != parent) {
spin_unlock(&dentry->d_lock);
node = rcu_dereference(node->next);
@@
if (dentry->d_name.len == name->len &&
memcmp(dentry->d_name.name, name->name, name->len) == 0) {
@@
if (!d_unhashed(dentry)) {
// Increment reference count before releasing the lock
atomic_inc(&dentry->d_count);
@@
node = rcu_dereference(node->next);
@@
return NULL;
Figure 14: Features implemented by SpecFS. For simplicity, a node in the figure may encompass multiple modules of specifications. “Common” denotes modifications made to some existing data structures, serving as the “rely” within other nodes’ corresponding specifications. “Root Identifier” indicates that this node is the root node, along with its associated logic.

### v3

@@ 6.1 Evaluations on Accuracy
Methodology. We first study whether SpecFS can accurately generate code to achieve self-evolution, that is, whether the code generation can accurately conform to the logic described by the specification and accurately meet the functionality of the corresponding file system module. Our test cases comprise all the modules to implement the complete logic of AtomFS [zou2019atomfs]. Specifically, to evaluate the accuracy, we first define 45 distinct modules within AtomFS and manually author their ground-truth implementations. A generated module is considered correct if it (i) passes all functional tests for AtomFS and (ii) is deemed logically equivalent to the ground-truth through manual inspection.
We implement two versions of baselines based on a few-shot learning approach. The normal version incorporates a description of the file correspondence logic and the APIs of the dependency modules; The oracle version not only includes the dependency module’s APIs but also integrates the ground-truth code of these modules as part of the prompt. We test generation accuracy using four LLMs of decreasing capability: Gemini-2.5-Pro [gemini], DeepSeek-V3.1 Reasoning [deepseekai2025deepseekv3technicalreport], GPT-5-minimal [gpt5], and Qwen3-32B [yang2025qwen3technicalreport], whose performance is ranked according to the LiveCodeBench leaderboard [livecodebench-leaderboard].
@@ 6.2 Evaluations on Generalizability
Generalizability for various locking methods. We evaluate whether the concurrency specifications of SysSpec are general for various locking methods. To this end, we utilize the dentry_lookup operation in the VFS layer of Linux, as it exhibits locking requirements at multiple granularities: a lock must be acquired for the entire hash list during traversal, while individual locks are also required for each dentry upon access. In our concurrency specification, we explicitly designate the use of two distinct locking mechanisms—lock-free RCU for the hash list and spinlocks for individual dentries. Our experiments verify that the generated code correctly adheres to the specified concurrency semantics, successfully producing code that implements multi-granularity and multi-method locking logic as intended. See details about the specification and the generated code of the two phases in the Appendix.
@@ 6.4 Evaluations on Productivity
Generation latency. The latency of code generation primarily depends on the inference performance of the LLM itself. In our experience, the generation time for SpecFS typically ranges from several minutes to tens of minutes.
@@ 6.6 Limitations and Discussion
Missing FS features. Although SpecFS represents a successful endeavor in realizing a complex file system using SysSpec, SpecFS is currently implemented as a user-space file system based on FUSE [fuse]. Consequently, it does not operate in kernel mode and lacks a storage stack, such as direct disk access, nor does it consider crash consistency. Our evaluation of SpecFS primarily focuses on validating the generation methodology and ensuring correctness, which precludes an apple-to-apple comparison with native kernel-space file systems regarding raw performance metrics such as throughput.
We plan to apply SpecFS to the self-evolution of industrial file systems (e.g., EROFS or Ext4 [gao2019erofs, linux-ext4]), which is challenging due to the increased engineering complexity. To address this, we consider a fully formally verified “AtomFS-Ext4” unnecessary. Instead, we propose developing a “SpecFS-Ext4” directly based on documentation, using these mitigation strategies. First, instead of specifying the whole state of Ext4, developers can start with a minimal baseline (like Ext2) and incrementally add features to it. Second, we plan to enhance the SpecAssistant to automatically bootstrap draft specifications from documentation (e.g., kernel wikis) or even Ext4 source code. Third, we could adapt methodologies of software engineering to guide the development of specification-constructed (rather than programming-language-constructed) file systems, e.g., applying methods similar to the Law of Demeter [demeterlaw] to reduce the coupling between modules.
@@ 7 Related Work
Other research addresses LLM-based code generation for specific domains outside file systems. QiMeng Xpiler [dong2025qimeng] uses LLMs to construct a transcompiler that adapts low-level tensor programs for various hardware platforms. Autoverus [yang2025autoverus] employs LLM-based tools for proving the correctness of Rust code. OSVBench and SpecGen [li2025osvbenchbenchmarkingllmsspecification, ma2025specgenautomatedgenerationformal] address generating formal specifications for a given code snippet.
Formal verification methods. Many studies verify complex systems, such as operating systems [gerwin2009sel4], file systems [zou2019atomfs, zou2024reffs, chen2016fscq], or cloud systems [sun2024anvil]. These efforts also involve formal specifications to ascertain the correctness of handcrafted implementations, but cannot be used diretly for generative file systems. We can draw upon these existing specifications when authoring SysSpec for code generation.
@@ 8 Conclusion
This paper presents SysSpec and SpecFS, a framework and a case for generative file systems. Different from traditional paradigms, SysSpec shifts the developer’s focus from writing low-level code to designing high-level specification, and SpecFS shows the potential benefits on evolvability.
Acknowledgments
We are grateful to our shepherd Mai Zheng for his detailed suggestions, which significantly improved the paper. We thank the anonymous FAST reviewers for their constructive feedback. This work was supported in part by the National Natural Science Foundation of China (No. 62432010, 62302300 and 62472279), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (JYB2025XDXM113), and the Fundamental Research Funds for the Central Universities. Corresponding authors: Dong Du (dd_nirvana@sjtu.edu.cn), Yubin Xia (xiayubin@sjtu.edu.cn), and Haibo Chen (haibochen@sjtu.edu.cn).
@@ A.1 Specification
struct hlist_head { /* */ };
struct hlist_node { /* */ };
@@ A.2 Code Snippet
Phase 1: Initial Implementation.
@@
unsigned int len = name->len;
const unsigned char *str = name->name;
[…]
@@
struct dentry *found = NULL;
[…]
struct dentry *dentry = hlist_entry(node, struct dentry, d_hash);
[…]
continue;
[…]
continue;
[…]
continue;
[…]
continue;
[…]
continue;
[…]
break;
[…]
found->d_count.counter++;
return found;
[…]
Phase 2: Refined Implementation.
struct hlist_node *node;
struct dentry *found = NULL;
unsigned int hash = name->hash;
@@
node = rcu_dereference(node->next)) {
@@
[…]
@@
memcmp(dentry->d_name.name, name->name, name->len) != 0) {
spin_unlock(&dentry->d_lock);
continue;
@@
spin_unlock(&dentry->d_lock);
@@
break;
[…]
Figure 14: Features implemented by SpecFS. For simplicity, a node in the figure may encompass multiple modules of specifications. “Root Identifier” indicates that this node is the root node, along with its associated logic.